OpenStudy (anonymous):

Can anyone help me put the equation x^2=4-7x into the quadratic equation formula, please:)

OpenStudy (amistre64):

might help to write it in ax^2 + bx + c = 0 format

OpenStudy (anonymous):

Yeah I know that its like the first step

OpenStudy (amistre64):

the quadratic formula is defined using the abc parts \[x=\frac{1}{2a}(-b\pm\sqrt{b^2-4ac})\]

OpenStudy (amistre64):

so, what do we have when it is in its ax^2 + bx + c = 0 format?

OpenStudy (anonymous):

Plug it in

OpenStudy (amistre64):

lets first see what you get when you work out the first step. unless of course youve got it from here

OpenStudy (anonymous):

No I don't that's kind of the reason I am asking for help

OpenStudy (amistre64):

this is the helping part :) we prefer that the askers participate in the solution process.

OpenStudy (anonymous):

Yeah I got that

OpenStudy (amistre64):

can you rearrange this into the ax^2 + bx + c = 0 x^2 = 4 -7x

OpenStudy (anonymous):

0=-x^2-7x+4

OpenStudy (amistre64):

thats pretty good

OpenStudy (anonymous):

So it's wrong? or no?

OpenStudy (amistre64):

its right enough :) i was expecting you to go the other way ... but its not wrong at all so lets compare ax^2 + bx + c = 0 -x^2 - 7x + 4 = 0 a= -1 b = -7 c = 4

OpenStudy (anonymous):

Now you plug in the variables,k so lets move on.

OpenStudy (amistre64):

\[x=\frac{1}{2a}(-b\pm\sqrt{b^2-4ac})\] \[x=\frac{1}{2(-1)}(-(-7)\pm\sqrt{(-7)^2-4(-1)(4)})\] \[x=-\frac{1}{2}(7\pm\sqrt{49+16})\] etc ...